# Agnostic calculation of atomic free energies with the descriptor density of states

### v1

@@ Abstract
We present a new method to evaluate vibrational free energies of atomic systems without a priori specification of an interatomic potential. Our model-agnostic approach leverages descriptors, high-dimensional feature vectors of atomic structure. The entropy of a high-dimensional density, the descriptor density of states, is accurately estimated with conditional score matching. Casting interatomic potentials into a form extensive in descriptor features, we show free energies emerge as the Legendre–Fenchel conjugate of the descriptor entropy, avoiding all high-dimensional integration. The score matching campaign requires less resources than fixed-model sampling and is highly parallel, reducing wall time to a few minutes, with tensor compression schemes allowing lightweight storage. Our model-agnostic estimator returns differentiable free energy predictions over a broad range of potential parameters in microseconds of CPU effort, allowing rapid forward and back propagation of potential variations through finite temperature simulations, long-desired for uncertainty quantification and inverse design. We test predictions against thermodynamic integration calculations over a broad range of models for BCC, FCC and A15 phases of W, Mo and Fe at high homologous temperatures. Predictions pass the stringent accuracy threshold of 1-2 meV/atom (1/40-1/20 kcal/mol) for phase prediction with propagated score uncertainties robustly bounding errors. We also demonstrate targeted fine-tuning, reducing the $\alpha-\gamma$ transition temperature in a non-magnetic machine learning model of Fe from 2030 K to 1063 K through back-propagation, with no additional sampling. Applications to liquids and fine-tuning foundational models are discussed along with the many problems in computational science which estimate high dimensional integrals.
@@
However, all current approaches still require a priori specification of MLIP parameters before any sampling is performed. The restriction to specific model parameters significantly complicates exploration of how MLIP parameters influence finite temperature properties, which has gained increasing interest for both uncertainty quantificationSwinburne and Perez (2025); Imbalzano et al. (2021); Musil et al. (2019); Best et al. (2024) in forward-propagation and an array of inverse design goals in back-propagationThaler et al. (2022); Lopanitsyna et al. (2023), such as targeted fine-tuningGrigorev et al. (2023); Maliyov et al. (2025).
@@
We discuss the many possible perspectives for our approach in the conclusions (section IX), including inverse design applications and extension to foundational MLIPsBatatia et al. (2023); Bochkarev et al. (2024). As our approach allows practical implementation of density of states methodsWang and Landau (2001); Pártay et al. (2021) to very high-dimensional systems, we anticipate application across the many areas of computational science that use linear feature modelsMarchand et al. (2023); Follain and Bach (2024); Ekborg-Tanner et al. (2024).
@@ II.3 Adiabatic Switching
In addition to the above methods which employ equilibrium averages, the adiabatic switchingAdjanor et al. (2006); de Koning et al. (1999); Freitas et al. (2016); Menon et al. (2024) method estimates free energy differences using the well-known Jarzynski equality Jarzynski (1997). The equality can be writtenFreitas et al. (2016)
@@ III.1 Admissible interatomic potentials
A wide variety of interatomic potentials can be cast into the general linear form (16). Clearly, these include the broad class of linear-in-descriptor (LML) models, including MTPShapeev (2016), ACFS Behler (2011, 2016), SNAPThompson et al. (2015), SOAP Bartók et al. (2013); Darby et al. (2022, 2023), ACELysogorskiy et al. (2021), MILADYGoryaeva et al. (2021a); Dézaphie et al. (2025) and PODNguyen (2023) descriptors, where
@@
LML models can reach extremely high ($<2$ meV/atom) accuracy to ab initio dataCastellano et al. (2024), with robust UQSwinburne and Perez (2025) and often excellent dynamical stability, essential for thermodynamic samplingZhu et al. (2017); Grabowski et al. (2019); Zhong et al. (2023); Zhu et al. (2024); Menon et al. (2024). Polynomial or kernel featurizations are regularly used to increase flexibility, e.g. qSNAPRohskopf et al. (2023), PiPAllen et al. (2021) GAPBartók and Csányi (2015); Deringer et al. (2019), n-body kernels Glielmo et al. (2017, 2018), kernelZhong et al. (2023) etc. For example, with qSNAP we have the quadratic featurization
@@
The generalized linear form can encompass more complex models if we only consider a subset of parameters adjustable, an approach adopted when fine-tuning recent message-passing neural network (MPNN) modelsBatatia et al. (2022, 2023); Bochkarev et al. (2024); Cheng (2024). In this case, we consider inputs of the MPNN readout layer $h({\mathbf{D}}_{i})$ as descriptors ${\mathbf{D}}_{i}$, then express the readout layer as a general linear model of the desired form (16). For example, in the MACE MPNN architectureBatatia et al. (2022), the readout layer is a sum of a linear function and a one layer neural network $f({\mathbf{D}}_{i})$, giving the featurization
@@ IV.2 The isosurface and descriptor entropies
The CD-DOS entropy $\mathcal{S}({\bm{\mathcal{D}}}|\alpha)$ measures the proportion of the isosurface phase space volume that has a global descriptor vector ${\bm{\mathcal{D}}}$, meaning descriptor values with larger $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ are more likely to be observed under unbiased isosurface sampling. Furthermore, $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ has two properties which greatly facilitate free energy estimation. Firstly, $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ is $N$-independent, i.e. intensive, as we prove in appendix D. Secondly, as $\Omega({\bm{\mathcal{D}}}|\alpha)$ is normalized, application of Laplace’s method (appendix A) gives a very useful result that fixes the maximum of $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$:
$\max_{{\bm{\mathcal{D}}}\in\mathbb{R}^{D}}\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)=0.$ (29)
@@ VI Score matching the conditional density of states
Evaluation of the free energy $\mathcal{F}_{{\mathbf{w}}}(\beta)$ in (37) requires a minimization over the intensive descriptor entropy $\mathcal{S}({\bm{\mathcal{D}}}|\alpha)$, equation (28). We will determine the parameters ${\bm{\Theta}}$ of a model $\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)$ through score matchingHyvärinen and Dayan (2005), to give estimates $\mathcal{F}_{{\mathbf{w}};{\bm{\Theta}}}(\beta)$. The estimated score reads
${\hat{\bm{\Psi}}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)}\equiv\nabla_{{\bm{\mathcal{D}}}}\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\in\mathbb{R}^{D},$ (44)
which has a score matching loss, using integration by partsHyvärinen and Dayan (2005) and $\langle\dots\rangle_{\alpha}$ for averages on isosurfaces $\hat{\alpha}({\mathbf{X}})=\alpha$,
$\mathcal{L}({\bm{\Theta}})\equiv\left\langle\frac{N}{2}\|{\hat{\bm{\Psi}}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)}\|^{2}+\nabla_{{\bm{\mathcal{D}}}}\cdot{\hat{\bm{\Psi}}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)}\right\rangle_{\alpha},$ (45)
where the factor of $N$ emerges as we consider the intensive log density rather than the log density. As discussed in appendix D, as $\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)$ is intensive moments of the score will scale asymptotically as $1/N$ and thus (45) is asymptotically intensive.
@@ VI.1 Low-rank compressed score models
While the developments of section V allow us to avoid high-dimensional integration, we still require a low-rank model to efficiently estimate and store any score model. In addition, the model should allow efficient minimization for free energy estimation via (37). Here, we use a common tensor compression approachSherman and Kolda (2020) to produce a low-rank score model for estimation of higher order moments.
@@
${\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)\equiv{\bf f}([{\bm{\mathcal{D}}}-{\bm{\mu}}_{\alpha}]\cdot{\bf v}_{\alpha,l})\in\mathbb{R}^{F},\quad l\in[1,D],$ (46)
giving a conditional entropy model, with ${{\bm{\Theta}}}_{l}(\alpha)\in\mathbb{R}^{F}$,
@@
The conditional score, equation (44), reads
${\hat{\bm{\Psi}}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)}=\sum_{l}\left(\partial{\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)\cdot{{\bm{\Theta}}}_{l}(\alpha)\right){\bf v}_{\alpha,l}\in\mathbb{R}^{F},$ (48)
@@
$\langle\partial{\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)[\partial{\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)]^{\top}\rangle_{\alpha}{{\bm{\Theta}}}_{l}(\alpha)\equiv\frac{-1}{N}\langle\partial^{2}{\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)\rangle_{\alpha}.$ (49)
@@ VII.2 Training data for Fe, W and Mo
The majority of the database configurations for Fe and W are those published in Goryaeva et al. (2021a). The W database originates from the defect- and dislocation-oriented database in Goryaeva et al. (2021a), which was modified and updated with molecular dynamics instances in Zhong et al. (2023) to improve its suitability for finite-temperature calculations and thermoelasticity of W. Finally, for this study, using the MLIP developed in Zhong et al. (2023), we prepared multiple samples of W in the A15 phase or liquid within the NPT ensemble, covering temperatures from 100 K to 5000 K. Each system contained 216 atoms. We selected 96 snapshots, which were then recomputed using the same DFT parameterization as in Goryaeva et al. (2021a); Zhong et al. (2023). The Mo database was specifically designed for this study to ensure a well-represented configuration of Mo at high temperatures in the bcc and A15 phases. The detailed components of the database, as well as the ab initio details, are described in the Appendix E.
@@ IX.2 Fine-tuning of foundational MLIPs
Section III.1 discussed the applicability of D-DOS beyond the linear MLIPs employed in our numerical experiments (section VII.1). In particular, the D-DOS scheme is well-suited to the fine-tuning of message passing neural network (MPNN) interatomic potentialsBatatia et al. (2022); Cheng (2024), which have gained prominence due to their ‘foundational’ ability to approximate broad regions of the periodic tableBatatia et al. (2023); Bochkarev et al. (2024). While MPNN training is a deep learning procedure involving millions of model weights, fine-tuning schemes typically only target a small number of parameters in the MPNN readout layer, which can be expressed in the general linear form (16). For example, uncertainty quantification schemes designed for linear models have recently been applied to e.g. the MACE-MPA-0 foundation modelPerez et al. (2025). The ability to fine-tune foundation models on available phase diagram data, either from experiment or from reference calculations, is actively sought in atomic modelling and will be investigated in the near future. In particular, the light computational demand and storage requirements of D-DOS mean production of a ‘foundational’ D-DOS sampling library for foundational MPNNs is feasible and indeed desirable, allowing practitioners to rapidly assess how e.g. fine-tuning on additional training data improves free energy predictions.
@@ Appendix D Intensivity of the descriptor entropy
where ${\bf v}\in\mathbb{R}^{D}$ and $J({\bf v}|\alpha)$ are clearly intensive. The interest of this cumulant generating function is to get the conditional density of states, and the descriptor entropy, using the inverse Laplace transform from this equality:
@@
It is possible to show the intensive nature of $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ through application of the inverse Laplace transform. However, as the cumulants of $\Omega({\bm{\mathcal{D}}}|\alpha)$ are finite, we know that $\Omega({\bm{\mathcal{D}}}|\alpha)$ has a global maximum at finite ${\bm{\mathcal{D}}}$. We can then apply Laplace’s method to (87) to find that, using the definition (28) of $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$,
$\lim_{N\to\infty}J({\bf v}|\alpha)=\max_{{\bm{\mathcal{D}}}}\mathcal{S}({\bm{\mathcal{D}}}|\alpha)+\mathbf{v}\cdot{\bm{\mathcal{D}}},\quad\Rightarrow\quad\mathcal{S}({\bm{\mathcal{D}}}^{*}_{\bf v}|\alpha)=J({\bf v}|\alpha)-\mathbf{v}\cdot{\bm{\mathcal{D}}}^{*}_{\bf v},$ (88)
where ${\bm{\mathcal{D}}}^{*}_{\bf v}$ solves the maximum condition.
As ${\bm{\mathcal{D}}}$ has a dimension $D$ that is independent of $N$, all prefactor terms vanish from the log integral as $N\to\infty$ (appendix A). We have thus established that $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ is intensive, as it is the sum of two manifestly intensive terms. As discussed in section V.2, the conditional free energy (36) can be expressed in terms of the cumulant generating function, with ${\bf v}=-\beta{\mathbf{w}}$; however, as we detail, we instead use score matching to estimate higher order moments.
Appendix E Databases design for Mo bcc and A15

### v3

@@ Abstract
We present a new method to evaluate vibrational free energies of atomic systems without a priori specification of an interatomic potential. Our model-agnostic approaches leverages descriptors, high-dimensional feature vectors of atomic structure. The entropy of a high-dimensional density, the descriptor density of states, is accurately estimated with conditional score matching. Casting interatomic potentials into a form extensive in descriptor features, we show free energies emerge as the Legendre–Fenchel conjugate of the descriptor entropy, avoiding all high-dimensional integration. The score matching campaign requires less resources than fixed-model sampling and is highly parallel, reducing wall time to a few minutes, with tensor compression schemes allowing lightweight storage. Our model-agnostic estimator returns differentiable free energy predictions over a broad range of potential parameters in microseconds of CPU effort, allowing rapid forward and back propagation of potential variations through finite temperature simulations, long-desired for uncertainty quantification and inverse design. We test predictions against thermodynamic integration calculations over a broad range of models for BCC, FCC and A15 phases of W, Mo and Fe at high homologous temperatures. Predictions pass the stringent accuracy threshold of 1-2 meV/atom (1/40-1/20 kcal/mol) for phase prediction with propagated score uncertainties robustly bounding errors. We also demonstrate targeted fine-tuning, reducing the $\alpha-\gamma$ transition temperature in a non-magnetic machine learning model of Fe from 2030 K to 1063 K through back-propagation, with no additional sampling. Applications to liquids and fine-tuning foundational models are discussed along with the many problems in computational science which estimate high dimensional integrals.
@@
However, all current approaches still require a priori specification of MLIP parameters before any sampling is performed. The restriction to specific model parameters significantly complicates exploration of how MLIP parameters influence finite temperature properties, which has gained increasing interest for both uncertainty quantificationSwinburne and Perez (2025); Imbalzano et al. (2021); Musil et al. (2019); Best et al. (2024); Zhu et al. (2023) in forward-propagation and an array of inverse design goals in back-propagationThaler et al. (2022); Lopanitsyna et al. (2023), such as targeted fine-tuningGrigorev et al. (2023); Maliyov et al. (2025).
@@
We discuss the many possible perspectives for our approach in the conclusions (section IX), including inverse design applications and extension to foundational MLIPsBatzner et al. (2022); Batatia et al. (2023); Bochkarev et al. (2024). As our approach allows practical implementation of density of states methodsWang and Landau (2001); Pártay et al. (2021) to very high-dimensional systems, we anticipate application across the many areas of computational science that use linear feature modelsMarchand et al. (2023); Follain and Bach (2024); Ekborg-Tanner et al. (2024).
@@ II.3 Adiabatic Switching
In addition to the above methods which employ equilibrium averages, the adiabatic switchingAdjanor et al. (2006); de Koning et al. (1999); Freitas et al. (2016); Menon et al. (2024) method estimates free energy differences using the well-known Jarzynski equality Jarzynski (1997). The adiabatic switching equality can be writtenFreitas et al. (2016)
@@ III.1 Admissible interatomic potentials
A wide variety of interatomic potentials can be cast into the general linear form (16). Clearly, these include the broad class of linear-in-descriptor (LML) models, including MTPShapeev (2016), ACFS Behler (2011, 2016), SNAPThompson et al. (2015), SOAP Bartók et al. (2013); Darby et al. (2022, 2023), ACEDrautz (2019, 2020); Lysogorskiy et al. (2021), MILADYGoryaeva et al. (2021a); Dézaphie et al. (2025) and PODNguyen (2023) descriptors, where
@@
LML models can reach extremely high ($<2$ meV/atom) accuracy to ab initio dataCastellano et al. (2024), with robust UQSwinburne and Perez (2025) and often excellent dynamical stability, essential for thermodynamic samplingZhu et al. (2017); Grabowski et al. (2019); Zhong et al. (2023); Zhu et al. (2024); Menon et al. (2024). Polynomial or kernel featurizations are regularly used to increase flexibility, e.g. qSNAPRohskopf et al. (2023), PiPAllen et al. (2021) GAPBartók and Csányi (2015); Deringer et al. (2019), n-body kernels Glielmo et al. (2017, 2018); Vandermause et al. (2020); Xie et al. (2021), kernelZhong et al. (2023) etc. For example, with qSNAP we have the quadratic featurization
@@
The generalized linear form can encompass more complex models if we only consider a subset of parameters adjustable, an approach adopted when fine-tuning recent message-passing neural network (MPNN) modelsBatzner et al. (2022); Musaelian et al. (2023); Batatia et al. (2022, 2023); Bochkarev et al. (2024); Cheng (2024); Batatia et al. (2025). In this case, we consider inputs of the MPNN readout layer $h({\mathbf{D}}_{i})$ as descriptors ${\mathbf{D}}_{i}$, then express the readout layer as a general linear model of the desired form (16). For example, in the MACE MPNN architectureBatatia et al. (2022), the readout layer is a sum of a linear function and a one layer neural network $f({\mathbf{D}}_{i})$, giving the featurization
@@ IV.2 The isosurface and descriptor entropies
The CD-DOS entropy $\mathcal{S}({\bm{\mathcal{D}}}|\alpha)$ measures the proportion of the isosurface phase space volume that has a global descriptor vector ${\bm{\mathcal{D}}}$, meaning descriptor values with larger $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ are more likely to be observed under unbiased isosurface sampling. Furthermore, $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ has two properties which greatly facilitate free energy estimation. In appendix D, we show that as the per-atom descriptor features ${\bm{\phi}}({\mathbf{D}}_{i})\in\mathbb{R}^{D}$ depend only on the local environment of atom $i$, the CD-DOS entropy $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ is intensive ($N$-independent). Secondly, as $\Omega({\bm{\mathcal{D}}}|\alpha)$ is normalized, application of Laplace’s method (appendix A) gives a very useful result that fixes the maximum of $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$:
$\lim_{N\to\infty}\frac{1}{N}\ln\left|\int\Omega({{\bm{\mathcal{D}}}}|\alpha){\rm d}{\bm{\mathcal{D}}}\right|=\max_{{\bm{\mathcal{D}}}\in\mathbb{R}^{D}}\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)=0.$ (29)
@@ VI Score matching the conditional density of states
Evaluation of the free energy $\mathcal{F}_{{\mathbf{w}}}(\beta)$ in (37) requires a minimization over the intensive descriptor entropy $\mathcal{S}({\bm{\mathcal{D}}}|\alpha)$, equation (28). We will determine the parameters ${\bm{\Theta}}$ of a model $\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)$ through score matchingHyvärinen and Dayan (2005), to give estimates $\mathcal{F}_{{\mathbf{w}};{\bm{\Theta}}}(\beta)$ with score $\bm{\nabla}\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\in\mathbb{R}^{D}$.
In appendix E we show the score matching lossHyvärinen and Dayan (2005) reads, using $\langle\dots\rangle_{\alpha}$ for averages on isosurfaces $\hat{\alpha}({\mathbf{X}})=\alpha$,
$\mathcal{L}({\bm{\Theta}}|\alpha)\equiv\left\langle\frac{N}{2}\|\bm{\nabla}\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\|^{2}+\bm{\nabla}^{2}\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\right\rangle_{\alpha}.$ (44)
The factor of $N$ emerges from application of integration by partsHyvärinen and Dayan (2005) in the derivation of (44). Although $\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)$ is intensive, averages over $\alpha$ of $\mathcal{S}({\bm{\mathcal{D}}}|\alpha)$ and it’s gradients will give rise to terms $\mathcal{O}(N^{-s})$, $s\geq 0$ in $\mathcal{L}({\bm{\Theta}}|\alpha)$, giving a multiscale hierarchyPavliotis and Stuart (2008) $\mathcal{L}({\bm{\Theta}}|\alpha)=\sum_{s}N^{1-s}\mathcal{L}_{s}({\bm{\Theta}}|\alpha)$. As we detail in appendix E, each $\mathcal{L}_{s}({\bm{\Theta}}|\alpha)$ should be minimized recursively, ensuring the solution at $s=S$ does not affect the solution at $s<S$. In practice, this gave negligable improvement over simply minimizing (44) directly, as for sufficiently large $N$ ($\mathcal{O}(100)$ in our experiments), a direct solution will naturally respect the dominant terms in the hierarchy, i.e. those for $s=0,1$.
@@ VI.1 Low-rank compressed score models
While the developments of section V allow us to avoid high-dimensional integration, we still require a low-rank model to efficiently estimate and store any score model. In addition, the model should allow efficient minimization for free energy estimation via (37). We use a common tensor compression approachSherman and Kolda (2020) to produce a low-rank model for estimation of higher order moments.
@@
${\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)\equiv{\bf f}([{\bm{\mathcal{D}}}-{\bm{\mu}}_{\alpha}]\cdot{\bf v}_{\alpha,l})\in\mathbb{R}^{F},\quad l\in[1,D].$ (45)
In practice, we use polynomial features of typical order $F=3-7$; we note that quadratic models ($F=2$) are insufficient to capture the anharmonic behavior of the D-DOS shown in section VIII. The conditional entropy model then reads, with ${{\bm{\Theta}}}_{l}(\alpha)\in\mathbb{R}^{F}$,
@@
The conditional descriptor score then reads
$\bm{\nabla}\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)=\sum_{l}\left(\partial{\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)\cdot{{\bm{\Theta}}}_{l}(\alpha)\right){\bf v}_{\alpha,l}\in\mathbb{R}^{F},$ (47)
@@
$N\langle\partial{\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)[\partial{\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)]^{\top}\rangle_{\alpha}{{\bm{\Theta}}}_{l}(\alpha)=-\langle\partial^{2}{\bf f}_{l}({\bm{\mathcal{D}}}|\alpha)\rangle_{\alpha}.$ (48)
@@ VII.2 Training data for Fe, W and Mo
The majority of the database configurations for Fe and W are those published in Goryaeva et al. (2021a). The W database originates from the defect- and dislocation-oriented database in Goryaeva et al. (2021a), which was modified and updated with molecular dynamics instances in Zhong et al. (2023) to improve its suitability for finite-temperature calculations and thermoelasticity of W. Finally, for this study, using the MLIP developed in Zhong et al. (2023), we prepared multiple samples of W in the A15 phase or liquid within the NPT ensemble, covering temperatures from 100 K to 5000 K. Each system contained 216 atoms. We selected 96 snapshots, which were then recomputed using the same DFT parameterization as in Goryaeva et al. (2021a); Zhong et al. (2023). The Mo database was specifically designed for this study to ensure a well-represented configuration of Mo at high temperatures in the bcc and A15 phases. The detailed components of the database, as well as the ab initio details, are described in the Appendix F.
@@ IX.2 Fine-tuning of foundational MLIPs
Section III.1 discussed the applicability of D-DOS beyond the linear MLIPs employed in our numerical experiments (section VII.1). In particular, the D-DOS scheme is well-suited to the fine-tuning of message passing neural network (MPNN) interatomic potentialsMusaelian et al. (2023); Batzner et al. (2022); Batatia et al. (2022); Cheng (2024), which have gained prominence due to their ‘foundational’ ability to approximate broad regions of the periodic tableBatatia et al. (2023); Bochkarev et al. (2024). While MPNN training is a deep learning procedure involving millions of model weights, fine-tuning schemes typically only target a small number of parameters in the MPNN readout layer, which can be expressed in the general linear form (16). For example, uncertainty quantification schemes designed for linear models have recently been applied to e.g. the MACE-MPA-0 foundation modelPerez et al. (2025). The ability to fine-tune foundation models on available phase diagram data, either from experiment or from reference calculations, is actively sought in atomic modelling and will be investigated in the near future. In particular, the light computational demand and storage requirements of D-DOS mean production of a ‘foundational’ D-DOS sampling library for foundational MPNNs is feasible and indeed desirable, allowing practitioners to rapidly assess how e.g. fine-tuning on additional training data improves free energy predictions.
@@ Appendix D Intensivity of the descriptor entropy
where ${\bf v}\in\mathbb{R}^{D}$ and $J({\bf v}|\alpha)$ are clearly intensive, meaning that we have the identity
@@
As the cumulants of $\Omega({\bm{\mathcal{D}}}|\alpha)$ are finite, we know that $\Omega({\bm{\mathcal{D}}}|\alpha)$ has a global maximum at finite ${\bm{\mathcal{D}}}$. As a result, we define the descriptor entropy as the (negative) Legendre-Fenchel transform of the cumulant generating function:
$\mathcal{S}({\bm{\mathcal{D}}}|\alpha)\equiv\min_{\bf v}J({\bf v}|\alpha)-{\bf v}\cdot{\bm{\mathcal{D}}}\equiv\lim_{N\to\infty}\frac{1}{N}\ln|\Omega({\bm{\mathcal{D}}}|\alpha)|,$ (87)
which is clearly both intensive and convex in ${\bm{\mathcal{D}}}$. We have thus established that $\mathcal{S}({{\bm{\mathcal{D}}}}|\alpha)$ is intensive, as it is the sum of two manifestly intensive terms. As discussed in section V.2, the conditional free energy (36) can be expressed in terms of the cumulant generating function, with ${\bf v}=-\beta{\mathbf{w}}$; however, as we detail, we instead use score matching to estimate higher order moments.
Appendix E Derivation of the score matching loss
Our starting point is the definition of the score matching loss, the Fisher divergenceHyvärinen and Dayan (2005)
$L({\bm{\Theta}}|\alpha)$ $\equiv\frac{N}{2}\langle\|\nabla\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)-\nabla\mathcal{S}({\bm{\mathcal{D}}}|\alpha)\|^{2}\rangle_{\alpha}-L_{0}$ (88)
$=\frac{N}{2}\langle\|\nabla\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\|^{2}\rangle_{\alpha}-\langle\nabla\mathcal{S}({\bm{\mathcal{D}}}|\alpha)\cdot\nabla\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\rangle_{\alpha},$ (89)
where $L_{0}$ is a ${\bm{\Theta}}$-independent constant and we set $L({\bm{\Theta}}|\alpha)$ to be $\mathcal{O}(N)$ by convention. As isosurface averages $\hat{\alpha}({\mathbf{X}})=\alpha$ can clearly be written an integral over the D-DOS $\Omega({\bm{\mathcal{D}}}|\alpha)=\exp[N\mathcal{S}({\bm{\mathcal{D}}}|\alpha)]$, we use integration by parts to write
$L({\bm{\Theta}}|\alpha)$ $\equiv\frac{N}{2}\langle\|\nabla\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\|^{2}\rangle_{\alpha}-N\int\nabla\mathcal{S}({\bm{\mathcal{D}}}|\alpha)\cdot\nabla\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\exp[N\mathcal{S}({\bm{\mathcal{D}}}|\alpha)]{\rm d}{\bm{\mathcal{D}}},$ (90)
$\equiv\frac{N}{2}\langle\|\nabla\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\|^{2}\rangle_{\alpha}+\langle\nabla\cdot\nabla\mathcal{S}_{\bm{\Theta}}({\bm{\mathcal{D}}}|\alpha)\rangle_{\alpha},$ (91)
as given in the main text.
E.1 Analysis in the limit $N\to\infty$
As we discuss in section VI, whilst the descriptor entropy $\mathcal{S}({\bm{\mathcal{D}}}|\alpha)$ is intensive, averages of $\mathcal{S}({\bm{\mathcal{D}}}|\alpha)$ or gradients over $\alpha$ will in general give rise to terms inversely proportional to $N$. To correctly infer the solution in the limit $N\to\infty$, we make the multiscale hierarchyPavliotis and Stuart (2008)
$L({\bm{\Theta}}|\alpha)=\sum_{s}N^{1-s}L_{s}({\bm{\Theta}}|\alpha)$ (92)
To solve this hierarchy we can in principle define a multiscale solution ${\bm{\Theta}}^{(S)}$ such that
$\nabla L({\bm{\Theta}}^{(S)}|\alpha)={\bf 0}+\mathcal{O}(N^{-S}).$ (93)
We can find a solution ${\bm{\Theta}}^{(S)}$ in a recursive fashion, first minimizing $L_{0}({\bm{\Theta}}|\alpha)$ to give ${\bm{\Theta}}^{(0)}$, then minimizing $L_{1}({\bm{\Theta}}|\alpha)$ under the constraint $\nabla L_{0}({\bm{\Theta}}|\alpha)=0$ to give ${\bm{\Theta}}^{(1)}$, and so on. However, we only consider models where $L({\bm{\Theta}}|\alpha)$ is linear in ${\bm{\Theta}}$, meaning the loss gradient can be decomposed as
$\sum_{s}N^{1-s}\left[{\bf A}_{s}{\bm{\Theta}}-{\bf b}_{s}\right]={\bf 0}.$ (94)
Respecting the multi-scale hierarchy then equates to ensuring ${\bm{\Theta}}^{(S+1)}-{\bm{\Theta}}^{(S)}$ is in the null space of all ${\bf A}_{s}$, $s\leq S$. Whilst we investigated solving each term in this hierarchy independently, in practice this had negligable improvement over simply minmizing the score matching loss (44) as the linear solve will naturally ensure the solution respects the multi-scale hierarchy to a within numerical tolerance.
Appendix F Databases design for Mo bcc and A15